# NeuroNAS: A Framework for Energy-Efficient Neuromorphic Compute-in-Memory Systems using Hardware-Aware Spiking Neural Architecture Search

### v2

@@ Abstract
Spiking Neural Networks (SNNs) have demonstrated capabilities for solving diverse machine learning tasks with ultra-low power/energy consumption. To maximize the performance and efficiency of SNN inference, the Compute-in-Memory (CIM) hardware accelerators with emerging device technologies (e.g., RRAM) have been employed. However, SNN architectures are typically developed without considering constraints from the application and the underlying CIM hardware, thereby hindering SNNs from reaching their full potential in accuracy and efficiency. To address this, we propose NeuroNAS, a novel framework for developing energy-efficient neuromorphic CIM systems using a hardware-aware spiking neural architecture search (NAS), i.e., by quickly finding an SNN architecture that offers high accuracy under the given constraints (e.g., memory, area, latency, and energy consumption). NeuroNAS employs the following key steps: (1) optimizing SNN operations to enable efficient NAS, (2) employing quantization to minimize the memory footprint, (3) developing an SNN architecture that facilitates an effective learning, and (4) devising a systematic hardware-aware search algorithm to meet the constraints. Compared to the state-of-the-art, NeuroNAS with 8bit weight precision quickly finds SNNs that maintain high accuracy by up to 6.6x search time speed-ups, while achieving up to 92% area savings, 1.2x latency speed-ups, 84% energy savings across CIFAR-10, CIFAR-100, and TinyImageNet-200 datasets; while the state-of-the-art fail to meet all constraints at once. In this manner, NeuroNAS enables efficient design automation in developing energy-efficient neuromorphic CIM systems for diverse ML-based applications.
Index Terms:
Spiking neural networks (SNNs), neuromorphic computing, compute-in-memory (CIM), neural architecture search (NAS)
@@ I Introduction
In recent years, neuromorphic computing has shown remarkable performance of spiking neural networks (SNNs) in solving diverse machine learning (ML) tasks with ultra-low-power/energy consumption. To maximize SNNs’ algorithmic performance and efficiency gains, many studies have been conducted, ranging from the developments of SNN models [1, 2, 3], hardware accelerators [4, 5, 6], as well as their deployments for diverse applications, such as image classification [7][8], object recognition in automobile [9][10], continual learning in autonomous systems [11, 12, 13, 14], embodied intelligence in robotics [15], and bio-signal processing in healthcare [16]. To maximize the efficiency of SNN processing, SNN hardware accelerators have been employed. However, they still suffer from high memory access energy that dominates the overall system energy (i.e., the memory wall bottleneck); see Fig. 1(a). To address this, the non-von Neumann Compute-in-Memory (CIM) paradigm with non-volatile memory (NVM) technologies, such as resistive random access memory (RRAM), has been investigated [17][18]. Specifically, CIM platforms minimize data movements between the memory and compute engine parts, hence the memory access energy.
@@
Figure 1: (a) Energy breakdown of SNN inference on von-Neumann-based neuromorphic accelerators: PEASE [19], TrueNorth [4], and SNNAP [20]; adapted from [21]. (b) Accuracy vs. memory of different SNNs on CIFAR-10: CIFARNet1 [22], CIFARNet2 [23], ResNet11 [24], ResNet19 [25], AutoSNN [1], and SpikeNAS [3]; adapted from studies in [1][3].
@@
However, simply executing the existing SNNs on a CIM platform may not fulfill the design requirements from the application (e.g., accuracy and latency) and hardware constraints (e.g., memory, area, and energy consumption), as SNN models are typically developed without considering the characteristics of the CIM hardware. Moreover, most of SNN models are derived from conventional Artificial Neural Networks (ANNs), whose operations are different from SNNs, thus they may lose unique SNN features (e.g., temporal information) that leads to sub-optimal accuracy [2]. This limitation is also observed in recent studies [2][3], i.e., designing SNNs in the spiking domain can achieve higher accuracy than deriving from the ANN domain; see Fig. 1(b). Furthermore, different ML-based applications have different requirements of algorithmic performance and efficiency. Therefore, to ensure the practicality of SNNs for real-world implementations, SNN developments have to meet the design requirements and constraints. In fact, maximizing SNN benefits for the targeted applications and CIM platforms requires SNN developments that can meet the expected performance, efficiency, and constraints, which is a non-trivial task. Moreover, manually developing the suitable SNN is time consuming and laborious, hence requiring an alternative technique that can find a desired solution efficiently.
@@
Research Problem: How to automatically and quickly develop an SNN architecture for neuromorphic CIM systems that achieves high accuracy and efficiency (for latency and energy consumption) under the given memory and area constraints? An efficient solution to this problem may enable automatic and quick developments of energy-efficient neuromorphic CIM systems for diverse ML applications.
I-A State-of-the-Art and Their Limitations
@@
Currently, state-of-the-art works focus on employing neural architecture search (NAS) for finding SNN architectures that achieve high accuracy [1, 2, 3][26, 27, 28, 29, 30]. However, all of these NAS methods do not consider multiple design requirements and constraints posed by the target application and the CIM hardware (i.e., memory, area, latency, and energy consumption), thereby limiting the applicability of neuromorphic CIM systems for diverse real-world applications. To illustrate the limitations of state-of-the-art and the optimization potentials, a case study is performed and discussed in Sec. I-B.
@@ I-B Case Study and Associated Research Challenges
We study the impact of applying constraints in NAS process. Here, we reproduce the state-of-the-art NAS for SNNs (i.e., SNASNet [2]), and then apply a memory constraint of 2x106 (2M) parameters on it to build a memory-aware SNASNet, while considering the CIFAR-100 dataset. For the processing hardware, we consider an RRAM-based CIM hardware platform from [18]. Note, the details of experimental setup are discussed in Sec. IV. The experimental results are shown in Fig. 2, from which we make the following key observations.
  • A smaller SNN may achieve comparable accuracy to the bigger one, showing possibility to obtain high accuracy with low memory.
  • Memory saving leads to the reduction of area, latency, and energy consumption of SNN processing on a CIM platform.
From these observations, we identify associated research challenges to be addressed for solving the research problem below.
  • The solution should consider multiple design requirements and constraints (i.e., memory, area, latency, and energy consumption) in its search algorithm to ensure the deployability of SNNs for the target application and the CIM platform.
@@
Figure 2: Experimental results of the case study considering the CIFAR-100: (a) number of weight parameters and accuracy; and (b) area, latency, and energy consumption when running networks on an RRAM-based CIM platform [18].
@@ I-C Our Novel Contributions
To address the research challenges, we propose NeuroNAS, a novel framework for quickly finding the SNN architecture that can achieve high accuracy, while meeting the application requirements and CIM constraints (i.e., memory, area, latency, and energy consumption) through hardware-aware spiking NAS. It is also the first work that develops an integrated hardware-aware spiking NAS for developing energy-efficient neuromorphic CIM systems. NeuroNAS employs the following key steps (see an overview in Fig. 3 and details in Fig. 5).
@@
    Optimization of SNN Operations: It aims to investigate and selects SNN operations that have positive impact on accuracy, while considering their memory costs. The selected operations are used for building a network architecture.
@@
    Development of Network Architectures: It aims to build a network that facilitates effective learning and efficient deployments on CIM. It leverages (1) the selection of SNN operations, and (2) the quantization-aware fitness evaluation policy to improve the quality of fitness function and optimize the network size.
@@
    A Hardware-aware Search Algorithm: It aims to find a network architecture that achieves high accuracy by searching for suitable neural architectures, while considering the given memory, area, latency, and energy constraints.
@@
Key Results: We evaluate NeuroNAS framework through a PyTorch implementation, and run it on Nvidia RTX A6000 GPU machines. Experimental results show that, NeuroNAS with 8bit weight precision quickly finds SNNs that maintain high accuracy by up to 6.6x search time speed-ups, while achieving up to 92% area savings, 1.2x latency speed-ups, 84% energy savings across CIFAR-10, CIFAR-100, and TinyImageNet-200 datasets compared to the state-of-the-art. Meanwhile, the state-of-the-art fail to meet all constraints at once.
@@ II Preliminaries
II-A NAS for SNNs
Overview: The focus of state-of-the-art works is still on the NAS development for finding SNN architectures that achieve high accuracy. These works can be categorized into two approaches. First, NAS with Training approach, which aims to train a super-network that contains all possible architecture candidates, and then search a subset of architecture that meets the fitness criteria; as shown in [1][27]. However, it incurs huge searching time, memory, and energy consumption. Second, NAS without Training approach, which overcomes the limitations of “NAS with Training” by directly evaluating the fitness scores of architecture candidates, then only training the one with the highest score [2][3][26][28, 29, 30]. However, all these state-of-the-art methods do not consider multiple design requirements and constraints posed by the target application and the CIM platform, hence limiting their applicability for diverse real-world applications.
NAS without Training: In this work, we consider the “NAS without training” approach due to its efficiency in finding an appropriate network without costly training. This concept was initially proposed for ANNs [31], then adopted in spiking domain. For instance, SNASNet [2] leverages the studies in [31] for evaluating the investigated SNN based on their representation capabilities, which indicate their potentials in learning effectively. It employs the following key steps.
@@ II-A NAS for SNNs
    It feeds mini-batch samples ($S$) to the network, then records the LIF neurons’ activities. If $U_{m}$ reaches $\theta$, the corresponding neuron is mapped to 1, and otherwise 0. In each layer, this mapping is represented as binary vector ($b$).
@@
    The Hamming distance $H(b_{i},b_{j})$ between samples $i$ and $j$ is calculated to construct a matrix ($\textbf{M}_{H}$); see Eq. 1. $N$ is the number of neurons in the investigated layer, while $\beta$ is the normalization factor to address high sparsity issue from LIF neurons.
@@
    Afterward, the fitness score ($F$) of each investigated architecture candidate is calculated using $F=\log\left(\det\;\lvert\sum_{l}\textbf{M}_{H}^{l}\rvert\right)$, and the one with highest-score is selected for training.
@@
In this work, we propose a similar approach but enhanced with quantization-aware mechanism in our NeuroNAS framework to adapt to the CIM configuration, which will be discussed in Sec. III-B2.
Neural Cell-based Strategy for NAS: It aims to provide a unified benchmark for NAS algorithms [32]. A neural cell (or simply cell) is a directed acyclic graph (DAG), that is originally designed with 4 nodes and 5 pre-defined operations, and its 2 nodes are connected to each other via a specific pre-defined operation [2][32]; see Fig. 4. Its idea is to search for the cell architecture and operations. In this work, we employ a similar neural cell-based strategy in our NeuroNAS framework but optimized for enabling a quick search and high accuracy, which will be discussed in Sec. III-A - Sec. III-B1.
Figure 4: The SNN macro-architecture with 2 neural cells, and each cell is a DAG whose edge denotes a specific pre-defined operation; adapted from [2].
Figure 5: Overview of our NeuroNAS framework, and its key steps for generating an SNN architecture that can achieve high accuracy while meeting the given memory, area, latency, and energy constraints; the novel contributions are highlighted in blue.
II-B CIM Hardware Architecture and SNN Mapping Strategy
Architecture: We employ the state-of-the-art CIM hardware for SNNs (i.e., SpikeFlow [18]), whose architecture is presented in Fig. 6. It comprises compute tiles (Tiles), a neuron module (NeurM), a global accumulator (GAccu), a pooling module (PoolM), and a global buffer (GBuff). These modules are connected to each other using network-on-chip (NoC) connection. Each tile comprises a tile input buffer (TIB), processing elements (PEs), a tile buffer (TBuff), and a tile accumulator (TAccu). Each PE comprises a PE input buffer (PIB), a number of crossbars (C), PE buffer (PBuff), and PE accumulator (PAccu). These modules are connected to each other using H-Tree connection. Meanwhile, each crossbar comprises input peripherals (IP), an array of NVM devices, multiplexers, analog-to-digital converter (ADC), shift-and-add circuit (SAC), and DIFF module to perform signed multiply-and-accumulate (MAC) operations.
Figure 6: The CIM hardware architecture that is considered in this work. It is based on the SpikeFlow architecture [18].
SNN Mapping: We employ the state-of-the-art mapping strategy on the CIM hardware in SpikeSim [18]; see Fig. 7. Its idea is to distribute weights across the crossbars, while exploiting weights and feature maps reuse. If we consider dimensions of $H$$\times$$H$$\times$$D$ for input feature maps (ifms), $P$$\times$$P$$\times$$D$$\times$$F$ for weights (wghs), and $X$$\times$$X$ for crossbar, then the mapping strategy can be described as follows.
  • Weight kernel elements in $P$$\times$$P$ are mapped on different crossbars. For instance, a $3$$\times$$3$ weight kernel needs 9 crossbars; see C1-C9.

### v3

# NeuroNAS: Enhancing Efficiency of Neuromorphic In-Memory Computing for Intelligent Mobile Agents through Hardware-Aware Spiking Neural Architecture Search

@@ Abstract
Intelligent mobile agents (e.g., UGVs and UAVs) typically demand low power/energy consumption when solving their machine learning (ML)-based tasks, since they are usually powered by portable batteries with limited capacity. A potential solution is employing neuromorphic computing with Spiking Neural Networks (SNNs), which leverages event-based computation to enable ultra-low power/energy ML algorithms. To maximize the performance efficiency of SNN inference, the In-Memory Computing (IMC)-based hardware accelerators with emerging device technologies (e.g., RRAM) can be employed. However, SNN models are typically developed without considering constraints from the application and the underlying IMC hardware, thereby hindering SNNs from reaching their full potential in performance and efficiency. To address this, we propose NeuroNAS, a novel framework for developing energy-efficient neuromorphic IMC for intelligent mobile agents using hardware-aware spiking neural architecture search (NAS), i.e., by quickly finding an SNN architecture that offers high accuracy under the given constraints (e.g., memory, area, latency, and energy consumption). Its key steps include: optimizing SNN operations to enable efficient NAS, employing quantization to minimize the memory footprint, developing an SNN architecture that facilitates an effective learning, and devising a systematic hardware-aware search algorithm to meet the constraints. Compared to the state-of-the-art techniques, NeuroNAS quickly finds SNN architectures (with 8bit weight precision) that maintain high accuracy by up to 6.6x search time speed-ups, while achieving up to 92% area savings, 1.2x latency improvements, 84% energy savings across different datasets (i.e., CIFAR-10, CIFAR-100, and TinyImageNet-200); while the state-of-the-art fail to meet all constraints at once. Therefore, NeuroNAS enables efficient design automation in developing neuromorphic IMC for energy-efficient mobile agents.
@@ I Introduction
Intelligent mobile agents (e.g., UGVs and UAVs) have shown potentials in helping humans in solving diverse, complex, and intensive machine learning (ML) tasks in more accurate and efficient way [1]. Apart from high accuracy, these agents require low power/energy execution of ML algorithms as they are usually powered by limited portable batteries; see Fig. 1(a). A potential solution is by employing neuromorphic computing with Spiking Neural Networks (SNNs) which can offer ultra-low power/energy ML algorithms through their very sparse event-based operations.
@@
To maximize the efficiency of SNN inference, specialized accelerators have been employed [2]. However, they suffer from high memory access energy which dominates the overall system energy (memory wall bottleneck); see Fig. 1(b). To address this, the non-von Neumann In-Memory Computing (IMC) paradigm with non-volatile memory (NVM) technologies, such as resistive random access memory (RRAM), has been investigated [3][4]. Specifically, IMC platforms minimize data movements between the memory and compute engine parts, hence the memory access energy; see Fig. 1(d).
@@
Figure 1: (a) Intelligent mobile agents are typically expected to provide high accuracy with low power/energy processing to solve ML-based tasks. (b) Breakdown of SNN inference energy on von-Neumann-based neuromorphic accelerators: PEASE [5], TrueNorth [6], and SNNAP [7]; adapted from [8]. (c) Accuracy vs. memory of different SNNs on CIFAR-10 dataset: CIFARNet1 [9], CIFARNet2 [10], ResNet11 [11], ResNet19 [12], AutoSNN [13], and SpikeNAS [14]; data are obtained from studies in [13][14]. (d) Energy efficiency of different types of single-core SNN accelerators: CMOS-based fully digital design, CMOS-based mixed-signal design, and RRAM-based IMC design; based on data from [2].
@@
However, simply executing the existing SNN models on an IMC platform may not fulfill the design requirements of mobile agents in terms of performance (e.g., accuracy and latency) and hardware constraints (e.g., memory, area, and energy consumption), as SNN models are usually developed without considering the characteristics of the IMC hardware. Moreover, most of SNN models are derived from conventional Artificial Neural Networks (ANNs), whose operations are different from SNNs, thus they may lose unique SNN features (e.g., temporal information) that leads to sub-optimal accuracy [15]. This limitation is also observed in recent studies [14][15], i.e., designing SNNs in the spiking domain can achieve higher accuracy than deriving SNNs from the ANN domain, as shown in Fig. 1(c). Furthermore, to ensure the practicality of SNNs for real-world implementation of mobile agents, SNN developments have to meet the application requirements and constraints. In fact, maximizing SNN benefits for mobile agent applications and IMC platforms requires SNN developments that can meet the expected performance, efficiency, and constraints, which is a non-trivial task. Moreover, manually developing the suitable SNN is time consuming and laborious, hence requiring an alternative technique that can find a desired solution efficiently.
@@
Research Problem: How to automatically and quickly develop an SNN architecture for neuromorphic-based IMC that achieves high accuracy and efficiency under the given application requirements and constraints? An efficient solution to this problem may enable automatic and quick developments of efficient neuromorphic IMC for intelligent mobile agents.
I-A State-of-the-Art of NAS for SNNs and Their Limitations
@@
Currently, state-of-the-art works focus on employing neural architecture search (NAS) for finding SNN architectures that achieve high accuracy [13, 15, 14, 16, 17, 18, 19, 20]. However, these NAS methods do not consider multiple design requirements and constraints posed by the targeted application and the IMC hardware (i.e., memory, area, latency, and energy consumption), thereby limiting the applicability of neuromorphic IMC for real-world mobile agents. To illustrate the limitations of state-of-the-art and the optimization potentials, a case study is performed and discussed in Sec. I-B.
@@ I-B Case Study and Associated Research Challenges
We study the impact of applying constraints in the NAS process. To do this, we reproduce the state-of-the-art NAS for SNNs (i.e., SNASNet [15]), and then apply a memory constraint of 2x106 (2M) weight parameters on it to build the memory-aware SNASNet, while considering the CIFAR-100 dataset. For the hardware, we consider an RRAM-based IMC platform from [4]. Note, details of experimental setup are discussed in Sec. IV. The experimental results are shown in Fig. 2, from which we make the following key observations.
Figure 2: Results from the case study considering the CIFAR-100 dataset: (a) number of weight parameters and accuracy; and (b) area, latency, and energy consumption when running SNNs on an RRAM-based IMC hardware [4].
  1. 1.
    A smaller SNN may achieve comparable accuracy to the larger one, and hence showing a possibility to obtain high accuracy with low memory.
  2. 2.
    Memory saving leads to the reduction of area, latency, and energy consumption of SNN processing.
From these observations, we identify research challenges to be addressed for solving the targeted problem, as follows.
  • The solution should consider multiple design requirements and constraints (i.e., memory, area, latency, and energy consumption) in its search algorithm to ensure the deployability of SNNs on the IMC platform for mobile agents.
@@ I-C Our Novel Contributions
To address the targeted problem and research challenges, we propose NeuroNAS, a novel framework for quickly developing the SNN architecture that can achieve high accuracy, while meeting the application requirements and IMC constraints (i.e., memory, area, latency, and energy consumption) through hardware-aware spiking NAS. It is also the first work that develops a hardware-aware spiking NAS for enhancing efficient neuromorphic IMC design targeting intelligent mobile agents. NeuroNAS employs the following key steps (see an overview in Fig. 3 and details in Fig. 7).
@@
    Optimization of SNN Operations (Sec. III-A): It investigates then selects SNN operations that have positive impact on the accuracy, while considering their memory costs. These operations will be used for building SNNs.
@@
    Development of Network Architectures (Sec. III-B): It aims to build an SNN architecture that facilitates effective learning and efficient deployments on IMC platform. It leverages the selected operations, and the quantization-aware fitness evaluation policy to improve the quality of fitness function and optimize the network size.
@@
    A Hardware-aware Search Algorithm (Sec. III-C): It aims to find the suitable SNN architecture that achieves high accuracy, while meeting the given requirements and constraints on memory, area, latency, and energy.
@@
Key Results: We evaluate NeuroNAS framework through a PyTorch implementation, and run it on Nvidia RTX A6000 Multi-GPU machines. As a result, NeuroNAS quickly finds SNNs (with 8bit weight precision) that maintain high accuracy by up to 6.6x search time speed-ups, while achieving up to 92% area savings, 1.2x latency improvements, 84% energy savings across CIFAR-10, CIFAR-100, and TinyImageNet-200 compared to the state-of-the-art. Meanwhile, the state-of-the-art fail to meet all constraints at once.
@@ II Preliminaries
II-A Spiking Neural Networks (SNNs)
SNNs draw inspiration from the biological neurons, i.e., utilizing spikes for conveying information and processing data [21, 22, 23]. Therefore, operations in SNNs depend on the neuronal dynamics of the spiking neuron model. Typically, a leaky integrate-and-fire (LIF) neuron model is used since it can provide diverse spike patterns with a low-complexity.
Figure 4: SNN macro-architecture with 2 neural cells, and each cell is a DAG whose edge denotes a specific pre-defined operation; adapted from [15].
Figure 5: The SNN IMC hardware architecture that is considered in this work. It is based on the SpikeFlow architecture [4].
Figure 6: SNN partitioning for mapping input feature maps (ifms) and weights (wghs) on the IMC crossbars; adapted from studies in [4]. The dimension of ifms is denoted as $H$$\times$$H$$\times$$D$, while wghs is denoted as $P$$\times$$P$$\times$$D$$\times$$F$.
Figure 7: Overview of our NeuroNAS framework, and its key steps for generating an SNN architecture that can achieve high accuracy while meeting the given memory, area, latency, and energy constraints from the mobile agent with neuromorphic IMC hardware; novel contributions are highlighted in blue.
II-B NAS for SNNs
Overview: State-of-the-art works still focus on the NAS development for finding SNN architectures that achieve high accuracy. They can be categorized into two approaches. First is “NAS with Training”, which trains a super-network that contains all possible architecture candidates, and then search a subset of architecture that meets the fitness criteria, as demonstrated in [13][17]. However, it incurs huge searching time, memory, and energy consumption. Second is “NAS without Training”, which directly evaluates the fitness scores of architecture candidates, and then only trains the one with the highest score [15, 14, 16][18, 19, 20]. However, all of these state-of-the-art methods have not considered multiple design requirements and constraints posed by the target application and the CIM platform, hence limiting their applicability for real-world mobile agent applications.
NAS without Training: In this work, we consider “NAS without training” approach [15, 14, 16][18, 19, 20] due to its efficiency in finding an appropriate network without costly training, as compared to “NAS with Training” approach [13][17]. This concept was initially proposed for ANNs [24], and then adopted in spiking domain. For instance, SNASNet [15] leverages the studies in [24] for evaluating the SNN based on its representation capabilities, which indicate the potentials in learning effectively. It employs the following key steps.
@@ II-B NAS for SNNs
    It feeds mini-batch samples ($S$) to network, then records the LIF neurons’ activities. If a neuron generates a spike, it is mapped to 1, and otherwise 0. In each layer, this mapping is represented as binary vector ($b$).
@@
    The Hamming distance $H(b_{i},b_{j})$ between samples $i$ and $j$ is calculated to construct a matrix ($\textbf{M}_{H}$); see Eq. 1. $N$ is the number of neurons in the investigated layer, while $\beta$ is the normalization factor to address high sparsity issue.
@@
    The fitness score ($F$) of each investigated architecture candidate is calculated using $F$=$\log\left(\det\;\lvert\sum_{l}\textbf{M}_{H}^{l}\rvert\right)$, and the one with highest-score is selected for training.
@@
In this work, we propose a similar approach but enhanced with quantization-aware mechanism in our NeuroNAS framework to adapt to the IMC hardware configuration, which will be discussed further in Sec. III-B2.
Neural Cell-based Strategy for NAS: It aims to provide a unified benchmark for NAS algorithms [25]. A neural cell (or simply cell) is a directed acyclic graph (DAG), that is originally designed with 4 nodes and 5 pre-defined operations, and its 2 nodes are connected to each other via a specific pre-defined operation [15][25], as shown in Fig. 4. Its idea is to search for the cell architecture and operations. In this work, we employ a similar neural cell-based strategy in our NeuroNAS framework but optimized for enabling a quick search and high accuracy, which will be discussed in Sec. III-A until Sec. III-B1.
II-C IMC Hardware Architecture and SNN Mapping Strategy
Architecture: We employ the state-of-the-art IMC hardware for SNNs (i.e., SpikeFlow [4]), whose architecture is shown in Fig. 5. It comprises multiple compute tiles (Tiles), a neuron module (NeurM), a global accumulator (GAccu), a pooling module (PoolM), and a global buffer (GBuff). These modules are connected to each other using network-on-chip (NoC). Each tile comprises a tile input buffer (TIB), processing elements (PEs), a tile buffer (TBuff), and a tile accumulator (TAccu). Each PE comprises a PE input buffer (PIB), a number of crossbars (C), PE buffer (PBuff), and PE accumulator (PAccu). These modules are connected to each other using H-Tree connection. Meanwhile, each crossbar comprises input peripherals (IP), an array of NVM devices, multiplexers, analog-to-digital converter (ADC), shift-and-add circuit (SAC), and DIFF module to perform signed multiply-and-accumulate (MAC) operations.
SNN Mapping: We employ the state-of-the-art mapping strategy on the IMC hardware from studies in [4]; see Fig. 6. Its idea is to distribute weights across the crossbars, while exploiting the weight- and feature map-reuse. If we consider dimensions of $H$$\times$$H$$\times$$D$ for input feature maps (ifms), $P$$\times$$P$$\times$$D$$\times$$F$ for weights (wghs), and $X$$\times$$X$ for crossbar, then the mapping strategy can be described as follows.
  • Weight kernel elements in $P$$\times$$P$ are mapped on different crossbars. For instance, a $3$$\times$$3$ weight kernel require 9 crossbars; see C1-C9.